what is the t-value for a 90% confidence interval with n = 10


1
Expert's answer
2021-06-04T04:18:28-0400

Based on the information provided, the significance level is "\alpha=0.1,"

"df=n-1=10-1=9" ​​degrees of fredom, and the critical value is "t_c=1.833113."